OpenStudy (anonymous):

I have to solve this equation by factoring it xsquared-64=0

OpenStudy (anonymous):

x=8 and -8

OpenStudy (anonymous):

(x+8)(x-8)=0 x=8 and -8

OpenStudy (anonymous):

thnx....can u help other prob please?

OpenStudy (anonymous):

yeah sure :-)

OpenStudy (anonymous):

4xsquared-25=0

OpenStudy (anonymous):



OpenStudy (anonymous):

(2x)^2 - 5^2=0 (2x-5)(2x+5)=0 x=(-5/2) and (5/2)

OpenStudy (anonymous):

is there a simpler way to solve that equation

OpenStudy (anonymous):

4x^2=25 x^2=(25/4) x+5/2 and -5/2

OpenStudy (anonymous):

how bout 100-xsquared=0